Crane Inc. uses a calendar year for financial reporting. The company is authorized to issue 8,210,000 shares of $12 par common stock. At no time has Crane issued any potentially dilutive securities. Listed below is a summary of Crane’s common stock activities. 1. Number of common shares issued and outstanding at December 31, 2015 2,040,000 2. Shares issued as a result of a 12% stock dividend on September 30, 2016 244,800 3. Shares issued for cash on March 31, 2017 1,870,000 Number of common shares issued and outstanding at December 31, 2017 4,154,800 4. A 2-for-1 stock split of Crane’s common stock took place on March 31, 2018 Compute the weighted-average number of common shares used in computing earnings per common share for 2016 on the 2017 comparative income statement. 2284800 shares

Compute the weighted-average number of common shares used in computing earnings per common share for 2017 on the 2017 comparative income statement. 3687300 shares

Compute the weighted-average number of common shares to be used in computing earnings per common share for 2017 on the 2018 comparative income statement. 7374600 shares

Compute the weighted-average number of common shares to be used in computing earnings per common share for 2018 on the 2018 comparative income statement. shares
Answer #1

Solution A:

Weighted average number of common shares to be used in computing earnings per common share for 2016 on the 2017 comparative income statement =

Outstanding shares at beginning of 2016 restated with stock dividend regardless of where in the year the stock dividends occurs

= 2040000*112% = 2284800 shares

Solution B:

Weighted average number of common shares to be used in computing earnings per common share for 2017 on the 2017 comparative income statement:

Outstanding shares at beginning of 2017 + Weighted average nos of shares issued during the year

= 2284800 + 1870000*9/12 = 3687300 shares

Solution C:

Weighted average number of common shares to be used in computing earnings per common share for 2017 on the 2018 comparative income statement.

= 2017 weighted average nos of share outstanding * retrospective adjustment of stock split

= 3687300 * 2 = 7374600 shares

Solution D:

Weighted average number of common shares to be used in computing earnings per common share for 2018 on the 2018 comparative income statement:

Nos of outstanding shares at beginning of 2018 * retrospective adjustment of stock split

= 4154800*2 = 8309600 shares